1. What is the empirical formulate for isopropyl alcohol (which contains only C, H and O)

if the combustion of a 0.255 grams isopropyl alcohol sample produces 0.561 grams of

CO2 and 0.306 grams of H2O?


2. Naphthalene, the active ingredient in one variety of mothballs, is an organic

compound that contains carbon and hydrogen only. Complete combustion of a

20.10 mg sample of naphthalene in oxygen yielded 69.00 mg of CO2 and 11.30 mg of

H2O. Determine the empirical formula of naphthalene.


3. Xylene, an organic compound that is a major component of many gasoline blends,

contains carbon and hydrogen only. Complete combustion of a 17.12 mg sample of

xylene in oxygen yielded 56.77 mg of CO2 and 14.53 mg of H2O. Determine the

empirical and molecular formula of xylene. Molecular weight of xylene is 106.16

g/mol.


1
Expert's answer
2021-02-24T03:47:02-0500

1. Mass of carbon in CO2= 12/44 x 0.561g = 0.153g

Mass of hydrogen in water = 2/18 x 0.306g = 0.034g

Mass of oxygen= 0.255-(0.153+0.034) = 0.068g

Let's find the mole ratios

C= 0.153/12= 0.0128

H= 0.034/1= 0.034

O=0.068/16= 0.00425

Divide by smallest

C= 0.0128/0.00425= 3

H= 0.034/0.00425 = 8

O= 0.00425/0.00425= 1

Empirical formula= C3H8O


2. Mass of C in CO2= 12/44 x 69.00 = 18.82mg

Mass of H in H2O= 2/18 x 11.30 = 1.256mg

Mole ratios

C= 18.82/12= 1.568

H= 1.256/1= 1.256

Divide by the smaller

C= 1.568/1.256= 1.2=1

H= 1.256/1.256= 1

E.F = CH